# Fundamental limitations on the recoverability of quantum processes

### v1

@@ C.6 Entropy gain under positive maps
For the given positive map $\mathcal{F}$, we define a positive and trace nonincreasing map $\widetilde{\mathcal{F}}$ as ~$\widetilde{\mathcal{F}}(\rho):=\frac{1}{\alpha}\mathcal{F}(\rho)$. Now we define an operator $\sigma:=\mathcal{F}(\rho)+\epsilon\mathbbm{1}$. Clearly for all $\epsilon>0$, we have $\sigma>0$. Let us consider the following:
@@
where in Eq. (138) we have used the definition of adjoint. As $\widetilde{\mathcal{F}}$ is trace nonincreasing we have ~*$\widetilde{\mathcal{F}}^{*}(\mathbbm{1})\leq\mathbbm{1}$ and hence we have used inequality (130) in the inequality (138). Now ~*$\widetilde{\mathcal{F}}^{*}(\sigma)=\widetilde{\mathcal{F}}^{*}\circ\mathcal{F}(\rho)+\epsilon\widetilde{\mathcal{F}}^{*}(\mathbbm{1})\leq\widetilde{\mathcal{F}}^{*}\circ\mathcal{F}(\rho)+\epsilon\mathbbm{1}$. Thus, using inequality (132) we have $D\left(\rho\;\middle\|\;\widetilde{\mathcal{F}}^{*}(\sigma)\right)\geq D\left(\rho\;\middle\|\;\widetilde{\mathcal{F}}^{*}\circ\mathcal{F}(\rho)+\epsilon\mathbbm{1}\right)$. Using this inequality, we have
$S(\sigma)-S(\rho)$ $\geq(1-\alpha)\operatorname{tr}(\rho\log\rho)+\alpha D\left(\rho\;\middle\|\;\widetilde{\mathcal{F}}^{*}\circ\mathcal{F}(\rho)\right)-\epsilon\operatorname{tr}(\log\sigma).$ (139)
@@
$=\operatorname{tr}(\rho\log\rho)-\operatorname{tr}(\rho\alpha\log\widetilde{\mathcal{F}}^{*}\circ\mathcal{F}(\rho))$
$=\operatorname{tr}(\rho\log\rho)-\operatorname{tr}(\rho\log\{\widetilde{\mathcal{F}}^{*}\circ\mathcal{F}(\rho)\}^{\alpha})$
$=D\left(\rho\;\middle\|\;\{\widetilde{\mathcal{F}}^{*}\circ\mathcal{F}(\rho)\}^{\alpha}\right)=D\left(\rho\;\middle\|\;\left(\frac{1}{\alpha}\right)^{\alpha}\left(\mathcal{F}^{*}\circ\mathcal{F}(\rho)\right)^{\alpha}\right).$
If we take $\mathcal{F}^{*}$ to be unital (or equivalently $\mathcal{F}$ to be a trace preserving), then we have $\alpha=1$, and we obtain a well-known result BDW (16): $S\left(\mathcal{F}(\rho)\right)-S(\rho)\geq D\left(\rho\;\middle\|\;\mathcal{F}^{*}\circ\mathcal{F}(\sigma)\right)$. ∎
@@ C.7 Proof of Entropy gain under superchannels
Let us consider a channel $\mathcal{N}:\mathcal{L}\left({A}\right)\rightarrow\mathcal{L}\left(B\right)$, such that the state that realizes the supremum of the entropy functional $S\left[\mathcal{N}\right]$ is an element of $\mathsf{FRank}\left(\widetilde{A}\otimes{A}\right)\cap\mathcal{D}\left(\widetilde{A}\otimes{A}\right)$, then we can write
@@
where |$\outerproduct{\Psi}{\Psi}\mathsf{FRank}\left(\widetilde{A}\otimes{A}\right)\cap\mathcal{D}\left(\widetilde{A}\otimes{A}\right)$. A simple calculation shows that $\left(\mathrm{id}_{\widetilde{A}}\otimes{R}_{A\rightarrow B}\right)\outerproduct{\Psi}{\Psi}=\Psi_{\widetilde{A}}\otimes\mathbbm{1}_{B}$, where $\Psi_{\widetilde{A}}=\mathrm{tr}_{A}\left(\outerproduct{\Psi}{\Psi}\right)$, and since $\Psi_{\widetilde{A}}$ is full rank operator, we have $\left(\mathrm{id}_{\widetilde{A}}\otimes\mathcal{N}_{A\rightarrow B}\right)\outerproduct{\Psi}{\Psi}={\mathsf{C}}_{\mathcal{N}}^{\Psi}$, where ${\mathsf{C}}_{\mathcal{N}}^{\Psi}$ is Choi operator of the quantum channel $\mathcal{N}_{A\rightarrow B}$ with respect to |$\outerproduct{\Psi}{\Psi}$ |. Thus, we can write the entropy of quantum channel $\mathcal{N}_{A\rightarrow B}$ as
@@
Now, consider a superchannel $\Theta:\mathcal{L}\left(A,B\right)\rightarrow\mathcal{L}\left(C,D\right)$ such that the state that realizes the supremum of the entropy functional $S\left[\Theta\left(\mathcal{N}_{A\rightarrow B}\right)\right]$ is an element of $\mathsf{FRank}\left(\widetilde{C}\otimes{C}\right)\cap\mathcal{D}\left(\widetilde{C}\otimes{C}\right)$. Again via a simple calculation we can show that $\left(\mathrm{id}_{\widetilde{C}}\otimes\Theta\left({R}_{A\rightarrow B}\right)\right)\outerproduct{\Phi}{\Phi}=\Phi_{\widetilde{C}}\otimes\mathbbm{1}_{D}$, where $\Phi_{\widetilde{C}}=\mathrm{tr}_{C}\left(\outerproduct{\Phi}{\Phi}\right)$, and since $\Phi_{\widetilde{C}}$ is full rank operator, we then have $\left(\mathrm{id}_{\widetilde{C}}\otimes\Theta\left(\mathcal{N}_{A\rightarrow B}\right)\right)\outerproduct{\Phi}{\Phi}=\mathsf{C}_{\Theta\left(\mathcal{N}\right)}^{\Phi}$, where $\mathsf{C}_{\Theta\left(\mathcal{N}\right)}^{\Phi}$ is Choi matrix for quantum channel $\Theta\left(\mathcal{N}_{A\rightarrow B}\right)$. Thus, we can write the entropy of quantum channel $\Theta\left(\mathcal{N}_{A\rightarrow B}\right)$ as
@@
So, the difference in the entropy of the quantum channel after and before applying the superchannel  comes out to be
@@
where $\Delta\left[\mathsf{C}_{\Theta(\mathcal{N})}^{\Phi},\mathsf{C}_{\mathcal{N}}^{\Psi}\right]:=S\left(\mathsf{C}_{\Theta\left(\mathcal{N}\right)}^{\Phi}\right)-S\left(\mathsf{C}_{\mathcal{N}}^{\Psi}\right)$ and $\Delta^{\prime}\left[\Psi_{\widetilde{A}},\Phi_{\widetilde{C}}\right]:=S\left(\Psi_{\widetilde{A}}\right)-S\left(\Phi_{\widetilde{C}}\right)$. We aim to find a lower bound for the entropy gain and one way to do this is to calculate a lower bound on either of the functions $\Delta$ or $\Delta^{\prime}$. One can also calculate a lower bound on $\Delta$ and $\Delta^{\prime}$ both. Now, one way to find a lower bound on $\Delta$ could be to prove the existence of a positive or unital(or sub-unital), and completely positive, and trace-preserving map, or completely positive map (see appendix C.6) which maps $\mathsf{C}^{\Phi}_{\Theta\left(\mathcal{N}\right)}$ to $\mathsf{C}_{\mathcal{N}}^{\Psi}$. Fortunately, in our case we do not have to prove the existence of a map, since we already have a completely positive map, called representing map, that maps $\mathsf{C}_{\mathcal{N}}^{\Psi}$ to $\mathsf{C}_{\Theta\left(\mathcal{N}\right)}^{\Phi}$, and for full rank Choi matrices, we denoted it with $\mathfrak{T}$, thus, we have $\mathsf{C}_{\Theta\left(\mathcal{N}\right)}^{\Phi}=\mathfrak{T}\left(\mathsf{C}_{\mathcal{N}}^{\Psi}\right)$. Now, we can find a lower bound on $\Delta\left[\mathsf{C}^{\Phi}_{\Theta\left(\mathcal{N}\right)},\mathsf{C}_{\mathcal{N}}^{\Psi}\right]$ in terms of relative entropy between Choi matrices $\mathsf{C}_{\mathcal{N}}^{\Psi}$ and $\left(\mathsf{C}_{\mathcal{N}}^{\Psi}\right)_{\alpha}$ using Theorem (3), which holds whenever $\mathfrak{T}$ is positive, as follows
@@
where $\left(\mathsf{C}_{\mathcal{N}}^{\Psi}\right)_{\alpha}=\frac{1}{\alpha^{\alpha}}\left(\mathfrak{T}^{*}\circ\mathfrak{T}\left(\mathsf{C}_{\mathcal{N}}^{\Psi}\right)\right)^{\alpha}=\frac{1}{\alpha^{\alpha}}\left(\mathsf{C}^{\Psi}_{{\hat{\mathrm{\Theta}}_{\Psi}}^{-1}\circ\mathrm{\Theta}^{*}\circ\hat{\mathrm{\Theta}}_{\Phi}\circ\mathrm{\Theta}(\mathcal{N})}\right)^{\alpha}$ and $\alpha=\norm{\mathfrak{T}^{*}(\mathbbm{1})}_{\infty}=\norm{\mathsf{C}^{\Psi}_{{\hat{\mathrm{\Theta}}_{\Psi}}^{-1}\circ\mathrm{\Theta}^{*}\circ\hat{\mathrm{\Theta}}_{\Phi}\circ\mathrm{\Theta}^{-1}_{\Phi}(\mathcal{R})}}_{\infty}$ with $\mathcal{R}$ being the completely depolarizing map in $\mathcal{L}\left(C,D\right)$. The resulting inequality provides a lower bound of $\Delta\left[\mathsf{C}_{\Theta(\mathcal{N})}^{\Phi},\mathsf{C}_{\mathcal{N}}^{\Psi}\right]$. If we assume that $\mathcal{N}$ is a quantum channel, $\Psi$ is a normalized state, and the representing map $\mathfrak{T}$ to be unital, then we have $D\left(\mathsf{C}_{\mathcal{N}}^{\Psi}\;\middle\|\;\left(\mathsf{C}_{\mathcal{N}}^{\Psi}\right)_{\alpha}\right)\geq 0$ by Klein’s inequality (we also have to assume $\alpha=1$). Thus, we can conclude that the following inequality will hold
@@
which also completes our proof.
@@ Appendix D Refinement of data processing inequality
Let us consider a pair of channels $\mathcal{N}:\mathcal{L}\left({A}\right)\rightarrow\mathcal{L}\left(A^{\prime}\right)$ and $\mathcal{M}:\mathcal{L}\left({A}\right)\rightarrow\mathcal{L}\left(A^{\prime}\right)$ such that the state that realizes the supremum in the relative entropy functional $D\left[\mathcal{N}\|\mathcal{M}\right]$ is an element of $\mathsf{FRank}\left(\widetilde{A}\otimes{A}\right)\cap\mathcal{D}\left(\widetilde{A}\otimes{A}\right)$, then we can write
@@
where |$\outerproduct{\Psi}{\Psi}\in\mathsf{FRank}\left(\widetilde{A}\otimes{A}\right)\cap\mathcal{D}\left(\widetilde{A}\otimes{A}\right)$, and $\mathrm{tr}_{A}\left(\outerproduct{\Psi}{\Psi}\right)=|\widetilde{A}|$ |, we then have $\left(\mathrm{id}_{\widetilde{A}}\otimes\mathcal{N}\right)\outerproduct{\Psi}{\Psi}={\mathsf{C}}_{\mathcal{N}}^{\Psi}$ and $\left(\mathrm{id}_{\widetilde{A}}\otimes\mathcal{N}\right)\outerproduct{\Psi}{\Psi}={\mathsf{C}}_{\mathcal{M}}^{\Psi}$, where ${\mathsf{C}}_{\mathcal{N}}^{\Psi}$ and ${\mathsf{C}}_{\mathcal{M}}^{\Psi}$ are Choi matrix for quantum channel $\mathcal{N}$ and $\mathcal{M}$, respectively. Thus, we can write the relative entropy between quantum channels $\mathcal{N}$ and $\mathcal{M}$ as
@@
Now, consider a superchannel $\Theta:\mathcal{L}\left(A,A^{\prime}\right)\rightarrow\mathcal{L}\left(C,C^{\prime}\right)$ such that the state that realizes the supremum in the relative entropy functional $D\left[\Theta(\mathcal{N})\;\middle\|\;\Theta(\mathcal{M})\right]$ is an element of $\mathsf{FRank}\left(\widetilde{C}\otimes{C}\right)\cap\mathcal{C}\left(\widetilde{A}\otimes{C}\right)$. In particular, let |$\ket{\Phi}$ realize the supremum of the relative entropy $D\left[\Theta(\mathcal{N})\;\middle\|\;\Theta(\mathcal{M})\right]$. Since $\mathrm{tr}_{C}\left(\outerproduct{\Phi}{\Phi}\right)=|\widetilde{C}|$ |, we then have $\left(\mathrm{id}_{\widetilde{C}}\otimes\Phi\left(\mathcal{N}\right)\right)\outerproduct{\Phi}{\Phi}=\mathsf{C}_{\Theta\left(\mathcal{N}\right)}^{\Phi}$ and $\left(\mathrm{id}_{\widetilde{C}}\otimes\Phi\left(\mathcal{M}\right)\right)\outerproduct{\Phi}{\Phi}=\mathsf{C}_{\Theta\left(\mathcal{M}\right)}^{\Phi}$, where $\mathsf{C}_{\Theta\left(\mathcal{N}\right)}^{\Phi}$ and $\mathsf{C}_{\Theta\left(\mathcal{M}\right)}^{\Phi}$ are Choi matrices of channels $\Theta\left(\mathcal{N}\right)$ and $\Theta\left(\mathcal{M}\right)$, respectively. Thus, we can write the relative entropy between quantum channels $\Theta(\mathcal{N})$ and $\Theta(\mathcal{M})$ as
@@
Also, from the data processing inequality, we have $D\left[\mathcal{N}\|\mathcal{M}\right]\geq D\left[\Theta\left(\mathcal{N}\right)\|\Theta\left(\mathcal{M}\right)\right]$, which also implies the following inequlity
@@
Since, the Choi operators $\mathsf{C}_{\mathcal{N}}^{\Psi}$, $\mathsf{C}_{\mathcal{M}}^{\Psi}$, $\mathsf{C}_{\Theta(\mathcal{N})}^{\Psi}$ and $\mathsf{C}_{\Theta(\mathcal{N})}^{\Psi}$ are connected via representing map $\mathfrak{T}$ of superchannel $\Theta$ as $\mathsf{C}_{\Theta\left(\mathcal{N}\right)}^{\Phi}=\mathfrak{T}\left(\mathsf{C}_{\mathcal{N}}^{\Psi}\right)$ and $\mathsf{C}_{\Theta\left(\mathcal{M}\right)}^{\Phi}=\mathfrak{T}\left(\mathsf{C}_{\mathcal{M}}^{\Psi}\right)$. Using this, we can rewrite the above entropic inequality as
@@
Recall that the representing map $\mathfrak{T}$ of superchannel $\Theta$ is, in general, a CP map that becomes trace-preserving if the adjoint of $\Theta$ becomes completely depolarising map preserving (see Lemma 6). But we can replace it with the following trace preserving map: $\mathfrak{T}^{\prime}(\mathsf{X}):=\mathfrak{T}(\mathsf{X})+\left[\mathrm{tr}(\mathsf{X})-\mathrm{tr}(\mathfrak{T}(\mathsf{X}))\right]\sigma_{0}$, where $\sigma_{0}\in\mathcal{D}(A\otimes A^{\prime})$ is some fixed state and $\mathsf{X}\in\mathcal{L}(C\otimes C^{\prime})$, such that we have $\mathfrak{T}^{\prime}(\mathsf{C}^{\Psi}_{\mathcal{N}})=\mathfrak{T}(\mathsf{C}^{\Psi}_{\mathcal{N}})$, and similarly for $\mathsf{C}^{\Psi}_{\mathcal{M}}$. Thus, we have
@@
Since, $\mathfrak{T}^{\prime}$ is a quantum channel and normalized Choi operators can be taken as quantum states, so, we can write the refined version of the above inequality using Theorem 2.1 of  ${}^{+}$ (18) as (take ${\mathsf{C}}_{\mathcal{N}}^{\Psi}=\rho,{\mathsf{C}}_{\mathcal{M}}^{\Psi}=\sigma$ and $\mathfrak{T}^{\prime}=\mathcal{N}$)
$D\left({\mathsf{C}}_{\mathcal{N}}^{\Psi}\;\middle\|\;{\mathsf{C}}_{\mathcal{M}}^{\Psi}\right)-D\left(\mathfrak{T}\left({\mathsf{C}}_{\mathcal{N}}^{\Psi}\right)\;\middle\|\;\mathfrak{T}\left(\mathsf{C}_{\mathcal{M}}^{\Psi}\right)\right)\geq-\log F\left(\mathsf{C}^{\Psi}_{\mathcal{N}},\left(\mathcal{R}\circ\mathfrak{T}^{\prime}\right)\mathsf{C}^{\Psi}_{\mathcal{N}}\right),$ (156)
thus by having
$D\left[\mathcal{N}\;\middle\|\;\mathcal{M}\right]-D\left[\Theta\left(\mathcal{N}\right)\;\middle\|\;\Theta\left(\mathcal{M}\right)\right]\geq-\log F\left(\mathsf{C}^{\Psi}_{\mathcal{N}},\left(\mathcal{R}\circ\mathfrak{T}^{\prime}\right)\mathsf{C}^{\Psi}_{\mathcal{N}}\right),$ (157)

### v2

@@ C.6 Entropy gain under positive maps
For the given positive map $\mathcal{F}$, we define a positive and trace nonincreasing map $\widetilde{\mathcal{F}}$ as ~$\widetilde{\mathcal{F}}(\rho):=\frac{1}{\alpha}\mathcal{F}(\rho)$, where $\alpha:=\norm{\mathcal{F}^{*}(\mathbbm{1}_{B})}_{\infty}$. Now we define an operator $\sigma:=\mathcal{F}(\rho)+\epsilon\mathbbm{1}_{B}$. Clearly for all $\epsilon>0$, we have $\sigma>0$. Let us consider the following:
@@
As $\widetilde{\mathcal{F}}$ is trace nonincreasing we have ~$\widetilde{\mathcal{F}}^{*}(\mathbbm{1}_{B})\leq\mathbbm{1}_{A}$, then the inequality in the fourth line of above relations follows from the inequality (138). Now ~$\widetilde{\mathcal{F}}^{*}(\sigma)=\widetilde{\mathcal{F}}^{*}\circ\mathcal{F}(\rho)+\epsilon\widetilde{\mathcal{F}}^{*}(\mathbbm{1}_{B})\leq\widetilde{\mathcal{F}}^{*}\circ\mathcal{F}(\rho)+\epsilon\mathbbm{1}_{A}$. Thus, using inequality (140), we have $D\left(\rho\;\middle\|\;\widetilde{\mathcal{F}}^{*}(\sigma)\right)\geq D\left(\rho\;\middle\|\;\widetilde{\mathcal{F}}^{*}\circ\mathcal{F}(\rho)+\epsilon\mathbbm{1}_{A}\right)$. Using this, we get
$S(\sigma)-S(\rho)$ $\geq(1-\alpha)\operatorname{tr}(\rho\log\rho)+\alpha D\left(\rho\;\middle\|\;\widetilde{\mathcal{F}}^{*}\circ\mathcal{F}(\rho)+\epsilon\mathbbm{1}_{A}\right)-\epsilon\operatorname{tr}(\log\sigma).$ (143)
@@
$=\operatorname{tr}(\rho\log\rho)-\operatorname{tr}\left(\rho\alpha\log\widetilde{\mathcal{F}}^{*}\circ\mathcal{F}(\rho)\right)$
$=\operatorname{tr}\left(\rho\log\rho\right)-\operatorname{tr}\left(\rho\log\left\{\widetilde{\mathcal{F}}^{*}\circ\mathcal{F}(\rho)\right\}^{\alpha}\right)$
$=D\left(\rho\;\middle\|\;\left\{\widetilde{\mathcal{F}}^{*}\circ\mathcal{F}(\rho)\right\}^{\alpha}\right)=D\left(\rho\;\middle\|\;\left(\frac{1}{\alpha}\right)^{\alpha}\left(\mathcal{F}^{*}\circ\mathcal{F}(\rho)\right)^{\alpha}\right).$
This completes the proof of the theorem.
If we take $\mathcal{F}^{*}$ to be unital (or equivalently $\mathcal{F}$ to be a trace preserving) in Theorem 3, then we have $\alpha=1$, and we obtain a well-known result BDW (16): $S\left(\mathcal{F}(\rho)\right)-S(\rho)\geq D\left(\rho\;\middle\|\;\mathcal{F}^{*}\circ\mathcal{F}(\sigma)\right)$.
@@ C.7 Proof of Entropy gain under superchannels
Let us consider a channel $\mathcal{N}:\mathcal{L}\left({A}\right)\rightarrow\mathcal{L}\left(B\right)$ such that the state that realizes the supremum in the entropy functional $S\left[\mathcal{N}\right]$ is an element of $\mathsf{FRank}\left(\widetilde{A}\otimes{A}\right)\cap\mathcal{D}\left(\widetilde{A}\otimes{A}\right)$. Then we can write
@@
where |$\outerproduct{\Psi}{\Psi}\in\mathsf{FRank}\left(\widetilde{A}\otimes{A}\right)\cap\mathcal{D}\left(\widetilde{A}\otimes{A}\right)$. A simple calculation shows that $\left(\mathrm{id}_{\widetilde{A}}\otimes{R}_{A\rightarrow B}\right)\outerproduct{\Psi}{\Psi}=\Psi_{\widetilde{A}}\otimes\mathbbm{1}_{B}$, where $\Psi_{\widetilde{A}}=\mathrm{tr}_{A}\left(\outerproduct{\Psi}{\Psi}\right)$. Since $\Psi_{\widetilde{A}}$ is full rank operator, we have $\left(\mathrm{id}_{\widetilde{A}}\otimes\mathcal{N}_{A\rightarrow B}\right)\outerproduct{\Psi}{\Psi}={\mathsf{C}}_{\mathcal{N}}^{\Psi}$, where ${\mathsf{C}}_{\mathcal{N}}^{\Psi}$ is a Choi state of the quantum channel $\mathcal{N}_{A\rightarrow B}$ with respect to |$\outerproduct{\Psi}{\Psi}$ |. Thus, we can write the entropy of quantum channel $\mathcal{N}_{A\rightarrow B}$ as
@@
Now, consider a superchannel $\Theta:\mathcal{L}\left(A,B\right)\rightarrow\mathcal{L}\left(C,D\right)$ such that the state that realizes the supremum of the entropy functional $S\left[\Theta\left(\mathcal{N}_{A\rightarrow B}\right)\right]$ is an element of $\mathsf{FRank}\left(\widetilde{C}\otimes{C}\right)\cap\mathcal{D}\left(\widetilde{C}\otimes{C}\right)$. Again via a simple calculation we can show that $\left(\mathrm{id}_{\widetilde{C}}\otimes\Theta\left({R}_{A\rightarrow B}\right)\right)\outerproduct{\Phi}{\Phi}=\Phi_{\widetilde{C}}\otimes\mathbbm{1}_{D}$, where $\Phi_{\widetilde{C}}=\mathrm{tr}_{C}\left(\outerproduct{\Phi}{\Phi}\right)$. Since $\Phi_{\widetilde{C}}$ is full rank operator, we then have $\left(\mathrm{id}_{\widetilde{C}}\otimes\Theta\left(\mathcal{N}_{A\rightarrow B}\right)\right)\outerproduct{\Phi}{\Phi}=\mathsf{C}_{\Theta\left(\mathcal{N}\right)}^{\Phi}$, where $\mathsf{C}_{\Theta\left(\mathcal{N}\right)}^{\Phi}$ is a Choi state for quantum channel $\Theta\left(\mathcal{N}_{A\rightarrow B}\right)$. Thus, we can write the entropy of quantum channel $\Theta\left(\mathcal{N}_{A\rightarrow B}\right)$ as
@@
Now using Eqs. (146) and (147), the difference in the entropy of the quantum channel after and before applying the superchannel becomes
@@
where $\Delta\left[\mathsf{C}_{\Theta(\mathcal{N})}^{\Phi},\mathsf{C}_{\mathcal{N}}^{\Psi}\right]:=S\left(\mathsf{C}_{\Theta\left(\mathcal{N}\right)}^{\Phi}\right)-S\left(\mathsf{C}_{\mathcal{N}}^{\Psi}\right)$ and $\Delta^{\prime}\left[\Psi_{\widetilde{A}},\Phi_{\widetilde{C}}\right]:=S\left(\Psi_{\widetilde{A}}\right)-S\left(\Phi_{\widetilde{C}}\right)$.
In the following, we aim to find a lower bound for the entropy gain. One may calculate a lower bound on the functions $\Delta$ and $\Delta^{\prime}$ separately or together. Let $\mathfrak{T}$ be the representing for superchannel $\Theta$. Since Choi states $\mathsf{C}_{\Theta\left(\mathcal{N}\right)}^{\Phi}$ and $\mathsf{C}_{\mathcal{N}}^{\Psi}$ are connected via the representing map $\mathfrak{T}$ (see Eq. (II) and Remark 3), we can find a lower bound on $\Delta\left[\mathsf{C}^{\Phi}_{\Theta\left(\mathcal{N}\right)},\mathsf{C}_{\mathcal{N}}^{\Psi}\right]$ in terms of relative entropy between Choi states $\mathsf{C}_{\mathcal{N}}^{\Psi}$ and $\left(\mathsf{C}_{\mathcal{N}}^{\Psi}\right)_{\alpha}$ using Theorem 3, as follows:
@@
where $\left(\mathsf{C}_{\mathcal{N}}^{\Psi}\right)_{\alpha}=\frac{1}{\alpha^{\alpha}}\left(\mathfrak{T}^{*}\circ\mathfrak{T}\left(\mathsf{C}_{\mathcal{N}}^{\Psi}\right)\right)^{\alpha}=\frac{1}{\alpha^{\alpha}}\left(\mathsf{C}^{\Psi}_{{\hat{\mathrm{\Theta}}_{\Psi}}^{-1}\circ\mathrm{\Theta}^{*}\circ\hat{\mathrm{\Theta}}_{\Phi}\circ\mathrm{\Theta}(\mathcal{N})}\right)^{\alpha}$ and $\alpha=\norm{\mathfrak{T}^{*}(\mathbbm{1})}_{\infty}=\norm{\mathsf{C}^{\Psi}_{{\hat{\mathrm{\Theta}}_{\Psi}}^{-1}\circ\mathrm{\Theta}^{*}\circ\hat{\mathrm{\Theta}}_{\Phi}\circ\mathrm{\Theta}^{-1}_{\Phi}(\mathcal{R})}}_{\infty}$ with $\mathcal{R}$ being the completely depolarizing map in $\mathcal{L}\left(C,D\right)$. The resulting inequality provides a lower bound of $\Delta\left[\mathsf{C}_{\Theta(\mathcal{N})}^{\Phi},\mathsf{C}_{\mathcal{N}}^{\Psi}\right]$. With this, we conclude that the following inequality holds
@@
which completes our proof.
@@ Appendix D Refinement of data processing inequality
Let us consider a pair of quantum channels $\mathcal{N}:\mathcal{L}\left({A}\right)\rightarrow\mathcal{L}\left(A^{\prime}\right)$ and $\mathcal{M}:\mathcal{L}\left({A}\right)\rightarrow\mathcal{L}\left(A^{\prime}\right)$ such that the state that realizes the supremum in the relative entropy functional $D\left[\mathcal{N}\|\mathcal{M}\right]$ is an element of $\mathsf{FRank}\left(\widetilde{A}\otimes{A}\right)\cap\mathcal{D}\left(\widetilde{A}\otimes{A}\right)$. Then we can write
@@
where |$\outerproduct{\Psi}{\Psi}\in\mathsf{FRank}\left(\widetilde{A}\otimes{A}\right)\cap\mathcal{D}\left(\widetilde{A}\otimes{A}\right)$, and $\mathrm{rank}(\mathrm{tr}_{A}\left(\outerproduct{\Psi}{\Psi}\right))=|\widetilde{A}|$ |. We then have $\left(\mathrm{id}_{\widetilde{A}}\otimes\mathcal{N}\right)\outerproduct{\Psi}{\Psi}={\mathsf{C}}_{\mathcal{N}}^{\Psi}$ and $\left(\mathrm{id}_{\widetilde{A}}\otimes\mathcal{N}\right)\outerproduct{\Psi}{\Psi}={\mathsf{C}}_{\mathcal{M}}^{\Psi}$, where ${\mathsf{C}}_{\mathcal{N}}^{\Psi}$ and ${\mathsf{C}}_{\mathcal{M}}^{\Psi}$ are Choi states for quantum channels $\mathcal{N}$ and $\mathcal{M}$, respectively. Therefore, in this case, we can write the relative entropy between quantum channels $\mathcal{N}$ and $\mathcal{M}$ as
@@
Now, consider a superchannel $\Theta:\mathcal{L}\left(A,A^{\prime}\right)\rightarrow\mathcal{L}\left(C,C^{\prime}\right)$ such that the state that realizes the supremum in the relative entropy functional $D\left[\Theta(\mathcal{N})\;\middle\|\;\Theta(\mathcal{M})\right]$ is an element of $\mathsf{FRank}\left(\widetilde{C}\otimes{C}\right)\cap\mathcal{D}\left(\widetilde{C}\otimes{C}\right)$. In particular, let |$\ket{\Phi}$ be the state that realizes the supremum of the relative entropy $D\left[\Theta(\mathcal{N})\;\middle\|\;\Theta(\mathcal{M})\right]$. Since $\mathrm{rank}\left(\mathrm{tr}_{C}\left(\outerproduct{\Phi}{\Phi}\right)\right)=|\widetilde{C}|$ |, we have $\left(\mathrm{id}_{\widetilde{C}}\otimes\Theta\left(\mathcal{N}\right)\right)\outerproduct{\Phi}{\Phi}=\mathsf{C}_{\Theta\left(\mathcal{N}\right)}^{\Phi}$ and $\left(\mathrm{id}_{\widetilde{C}}\otimes\Theta\left(\mathcal{M}\right)\right)\outerproduct{\Phi}{\Phi}=\mathsf{C}_{\Theta\left(\mathcal{M}\right)}^{\Phi}$, where $\mathsf{C}_{\Theta\left(\mathcal{N}\right)}^{\Phi}$ and $\mathsf{C}_{\Theta\left(\mathcal{M}\right)}^{\Phi}$ are Choi states of channels $\Theta\left(\mathcal{N}\right)$ and $\Theta\left(\mathcal{M}\right)$, respectively. Thus, again, we can write the relative entropy between quantum channels $\Theta(\mathcal{N})$ and $\Theta(\mathcal{M})$ as
@@
Also, from the data processing inequality, we have $D\left[\mathcal{N}\|\mathcal{M}\right]\geq D\left[\Theta\left(\mathcal{N}\right)\|\Theta\left(\mathcal{M}\right)\right]$, which then implies the following inequality
@@
Since, the Choi states $\mathsf{C}_{\mathcal{N}}^{\Psi}$, $\mathsf{C}_{\mathcal{M}}^{\Psi}$, $\mathsf{C}_{\Theta(\mathcal{N})}^{\Phi}$ and $\mathsf{C}_{\Theta(\mathcal{N})}^{\Phi}$ are connected via representing map $\mathfrak{T}$ of superchannel $\Theta$ as $\mathsf{C}_{\Theta\left(\mathcal{N}\right)}^{\Phi}=\mathfrak{T}\left(\mathsf{C}_{\mathcal{N}}^{\Psi}\right)$ and $\mathsf{C}_{\Theta\left(\mathcal{M}\right)}^{\Phi}=\mathfrak{T}\left(\mathsf{C}_{\mathcal{M}}^{\Psi}\right)$. Using this, we can rewrite the above entropic inequality as
@@
Recall that the representing map $\mathfrak{T}$ of superchannel $\Theta$ is, in general, a CP map that becomes trace preserving if the adjoint of $\Theta$ becomes completely depolarising map preserving (see Lemma 6). But, in general, we can replace $\mathfrak{T}$ with the following trace preserving map: $\mathfrak{T}^{\prime}(\mathsf{X}):=\mathfrak{T}(\mathsf{X})+\left[\mathrm{tr}(\mathsf{X})-\mathrm{tr}(\mathfrak{T}(\mathsf{X}))\right]\sigma_{0}$, where $\mathsf{X}\in\mathcal{L}(C\otimes C^{\prime})$ and $\sigma_{0}\in\mathcal{L}(A\otimes A^{\prime})$ is chosen such that $\mathfrak{T}^{\prime}$ is a CPTP map, which further satisfy $\mathfrak{T}^{\prime}(\mathsf{C}^{\Psi}_{\mathcal{N}})=\mathfrak{T}(\mathsf{C}^{\Psi}_{\mathcal{N}})$ and $\mathsf{C}^{\Psi}_{\mathcal{M}}$ (see Appendix B.4). This is possible as $\Theta\in\mathsf{SCT}$. Thus, we have
@@
Since $\mathfrak{T}^{\prime}$ is a quantum channel and Choi states are quantum states, we can write the refined version of the above inequality using the result (JRS+, 18, Theorem 2.1) as
$D\left({\mathsf{C}}_{\mathcal{N}}^{\Psi}\;\middle\|\;{\mathsf{C}}_{\mathcal{M}}^{\Psi}\right)-D\left(\mathfrak{T}\left({\mathsf{C}}_{\mathcal{N}}^{\Psi}\right)\;\middle\|\;\mathfrak{T}\left(\mathsf{C}_{\mathcal{M}}^{\Psi}\right)\right)\geq-\log F\left(\mathsf{C}^{\Psi}_{\mathcal{N}},\left(\mathcal{P}^{\mathrm{R}}\circ\mathfrak{T}^{\prime}\right)\mathsf{C}^{\Psi}_{\mathcal{N}}\right).$ (157)
Therefore, we have
$D\left[\mathcal{N}\;\middle\|\;\mathcal{M}\right]-D\left[\Theta\left(\mathcal{N}\right)\;\middle\|\;\Theta\left(\mathcal{M}\right)\right]\geq-\log F\left(\mathsf{C}^{\Psi}_{\mathcal{N}},\left(\mathcal{P}^{\mathrm{R}}\circ\mathfrak{T}^{\prime}\right)\mathsf{C}^{\Psi}_{\mathcal{N}}\right),$ (158)